Ahmed & Co. makes and sells two types of shoes, Plain and Fancy. Data concerning these products are as follows:

Plain Fancy
Unit selling price $ 25.00 $ 36.00
Variable cost per unit 13.00 20.00

Sixty percent of the unit sales are Plain, and annual fixed expenses are $95,200.

Assuming that the sales mix remains constant, the total number of units that Ahmed must sell to break even is: (Round intermediate calculations to 2 decimal places and final answer to nearest whole number)
Answer #1

Weighted average contribution margin per unit = (25-13*60%)+(36-20*40%) = 13.60

Break even Unit = Fixed cost/Weighted average contribution margin per unit = 95200/13.6 = 7000 Units